OpenStudy (anonymous):

Find solution of quadratic equation: 5x^2=10x

sam (.sam.):

divdide both sides by 5 5x^2=10x x^2=2x divide both sides by x and you're done

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 5x^2 - 10x = 0 }\) \(\Large \color{MidnightBlue}{\Rightarrow 5x(x - 2) = 0 }\) \(\Large \color{MidnightBlue}{\Rightarrow x - 2 = 0 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 2 }\)

OpenStudy (anonymous):

The answer key tells me the answer is x=0, 2...where is the 0 coming from?

OpenStudy (amorfide):

parthkohli is wrong 5x²-10x=0 5x(x-2)=0 x=2 5x=0 x=0 x=2 , x=0

OpenStudy (anonymous):

Thanks :D

OpenStudy (amorfide):

anytime, happy to help :D

OpenStudy (amorfide):

if you follow sams method and divide by x you eliminate an answer you eliminate the possibility of x=0 don't follow that method you will only get half an answer